OpenStudy (gabylovesyou):

LAST QUESTION!!Part 1: Solve the inequality 12 + 4x

OpenStudy (gabylovesyou):

@jim_thompson5910

OpenStudy (anonymous):

have you solved for x yet?

jimthompson5910 (jim_thompson5910):

go ahead and solve for x and tell me what you get you don't have to completely solve if you get stuck

jimthompson5910 (jim_thompson5910):

just show us what you have so far

OpenStudy (gabylovesyou):

x >/= 14

OpenStudy (anonymous):

great !

OpenStudy (gabylovesyou):

i just have to put it step by step for part one but idk how to lol

OpenStudy (anonymous):

written explanation? or actually showing the work?

OpenStudy (gabylovesyou):

showing the work

jimthompson5910 (jim_thompson5910):

12 + 4x <= 5x - 2 12 <= 5x - 2 - 4x 12 <= x - 2 12+2 <= x - 2+2 14 <= x x >= 14

jimthompson5910 (jim_thompson5910):

That's what you should have. Of course, there's more than one way to do this.

jimthompson5910 (jim_thompson5910):

Now just graph it or describe how to graph it

OpenStudy (gabylovesyou):

uhhhh.............. well i know its a closed circle right??

jimthompson5910 (jim_thompson5910):

it is

jimthompson5910 (jim_thompson5910):

we're including the endpoint

OpenStudy (gabylovesyou):

and its shading to the left??

jimthompson5910 (jim_thompson5910):

>= is greater than or equal to so x >= 14 means "any number larger than 14 or 14 itself"

OpenStudy (gabylovesyou):

How would i put Part 2? How would i explain how to graph it?

jimthompson5910 (jim_thompson5910):

Just describe the endpoint (open or closed), then describe the shading (left or right)

jimthompson5910 (jim_thompson5910):

that should be a good substitute for the actual graph

OpenStudy (gabylovesyou):

the endpoint is closed and the shading is to the left right?

jimthompson5910 (jim_thompson5910):

why to the left?

OpenStudy (gabylovesyou):

cuz it shades wherever the less than sign is right???

jimthompson5910 (jim_thompson5910):

not sure what you mean, but... if you have < or <=, you shade to the left if you have > or >=, you shade to the right

OpenStudy (gabylovesyou):

ohhhhhhh ok

OpenStudy (gabylovesyou):

can you put it together??

jimthompson5910 (jim_thompson5910):

the shading to the right represents all the numbers that are greater than or equal than 14

OpenStudy (gabylovesyou):

Can you put it all together like if you were answering part 2 ??

jimthompson5910 (jim_thompson5910):

To graph x >= 14, draw a number line. Then plot a closed circle at 14. Then shade to the right of this closed circle.

OpenStudy (gabylovesyou):

Thank you sooo much!! :)

jimthompson5910 (jim_thompson5910):

np